OpenStudy (anonymous):

7(x -2) + 3(x + 2) = 5(x - 6)

OpenStudy (saifoo.khan):

7x - 14 + 3x + 6 = 5x - 30

OpenStudy (aroub):

how amny questions left?!?!

OpenStudy (anonymous):

addd man

OpenStudy (saifoo.khan):

5x = 22 x = 22/5

OpenStudy (anonymous):

7x-14+3x+6=5x-30 5x=-22 x=-22/5

OpenStudy (anonymous):

yeah am right

OpenStudy (anonymous):

saifo ur wrong